OpenStudy (anonymous):

Find the greatest common factor of 270 and 360.

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

@MrNood

OpenStudy (mrnood):

Starting at 360 write the factors: 360, 180, 120, 90 .... etc Do the same for 270 There is 1 obvious common factor which is the highest which is in BOTH lists

OpenStudy (anonymous):

idek the 2nd list

OpenStudy (mrnood):

That's cos oyu have to do it 360/1 = 360 360/2= 180 360/3 = 120 etc 270/1= 270 270/2= 135 If the result is not a whole number then it is NOT a factor

OpenStudy (anonymous):

i still dont get it can you give me the answer

OpenStudy (mrnood):

No way Do you know what a factor is? If so then write down the factors as I described above if not then go and find out

OpenStudy (anonymous):

ALRIGHT ill try this

OpenStudy (anonymous):

look at the question in the picture ik 3^2 is an answer

OpenStudy (anonymous):

@MrNood

OpenStudy (anonymous):

@AmourKiaraa

OpenStudy (mrnood):

It still seems like you are waiting for me to give you an answer My method which I described above will give you a VERY quick answer to this particular question because this is a very simple example. However - your answer sheet required you to do the prime factor method, which works for more complicated cases and it is best if you use this and UNDERSTAND it I have no interest in giving you the answer - that merely proves that I can do it. Please use the method you have been taught and make some effort to do it yourself. Once you have demonstrated some positive effort I may be able to help further.